OpenStudy (lovelyanna):

Which of the following has a slope of –2 and a y-intercept of 4? A. y = 2x + 4 B. y = –2x – 4 C. y = –2x + 4 D. y = 2x – 4 Is it D?

OpenStudy (lovelyanna):

@shamil98

OpenStudy (shamil98):

y = mx +b slope = m y intercept = b you got the positive signs mixed up, but good try!

OpenStudy (lovelyanna):

So this would be B?

OpenStudy (shamil98):

No, the y intercept is positive 4 you got the slope right this time though

OpenStudy (lovelyanna):

Oh.. A?

OpenStudy (shamil98):

No. It is C. y = -2x + 4 y = mx + b m = -2 y = 4

OpenStudy (lovelyanna):

Oh.. sorry :(

OpenStudy (shamil98):

good try, though. just remember you don't need to look at the answer choices you are given m and b just plug those into y = mx + b y = (-2)x + (4) y = -2x + 4

OpenStudy (lovelyanna):

Oh, ok. Thanks!